Yesterday Alex had a new temp. central line place. After the procedure was compleated, he was able to have his scalp IV removed, however, once they removed it he bled a good bit. A little while late his new line began leaking some blood. The nurses cleaned it and put a new dressing on it but it was still leaking just a little. When they took his labs in the afternoon, his H&H came back very low. He received his third blood transfusion from around midnight until 4am. He had a good day today otherwise. He did have some minor issues with one side of his new line clotting off. The nurse put a drug called TPA into his line and let it sit for 1.5 hours and then was able to pull the clots out. There is still no news on the bed situation in CHOP. The hope is that we will going tomorrow mid day but who knows at this point!!